[Sara-Reports] SARA KPNO STATUS: BACK ONLINE

Oswalt, Terry D. oswaltt1 at erau.edu
Thu Dec 15 07:35:09 CST 2016


Thanks, Peter.
You have my permission to call out an electrician to check the building wiring, if you think it is warranted.
Perhaps a mouse has gnawed the insulation off a wire…

Terry D. Oswalt, Ph.D.
Chair, Dept. Physical Sciences
Engineering Physics – Space  Physics – Astronomy & Astrophysics
Embry-Riddle Aeronautical University
600 Clyde Morris Boulevard
Daytona Beach, FL 32114
Office: (386) 226-7571
Fax: (386) 226-6621
Terry.Oswalt at erau.edu<mailto:Terry.Oswalt at erau.edu>
[cid:image001.png at 01D256AE.24B16730]

From: <sara-reports-bounces at tamuc.edu> on behalf of "pmack at astronomical.com" <pmack at astronomical.com>
Date: Wednesday, December 14, 2016 at 7:31 PM
To: "sara-reports at tamuc.edu" <sara-reports at tamuc.edu>
Subject: [Sara-Reports] SARA KPNO STATUS: BACK ONLINE

James and I visited the mountain today.

CAMERA:  Now back in operation, cooling very nicely.  Only took 1-1/2 hours to get to -107C.

COOLER:  I confirmed Tony’s observation that the old cooler is in need of maintenance, beyond what we can do.
We have brought it back to Tucson for shipment back to the factory.

WIRING: So, something strange was going on.  The 20 Amp breaker that deals with the outlets around the base of the dome was tripping.
Courtesy of the cooler.  However, the outlet that the cooler was originally plugged into was not working correctly.  It has only (approx) 30 VAC on the outlet.
All the other outlets are working normally.  The electric roller shutter door was also wired into the back of the outlet.  I rewired the door so it is on the UPS.
So now we have a working door again.  And I have isolated the part of the wiring giving 30 VAC. So now we have a puzzle.  Did the cooler cause a fault in the building wiring.  Or did the building wiring cause a fault on the cooler?  We know that they must be related (both failed at the same time).

LOCKS:  The lock for the lower door has failed.  On the next visit we will put the upstairs lock on the downstairs door (hardly ever used) and put a new lock on the main door.  (About $160).

STELLA CAM:  It was working all the time last week.  Apparently it was put on camera channel 6, on the second screen, so it was not obvious.  I have moved it back to channel 4.
However, we are still getting a new detector to try out.  More on that as we do some bench tests in Tucson.

TELESCOPE WORM:  I found that the telescope was out of balance in HA, being too heavy on the east (Tube) side.  I adjusted it with the rolling weights, so it is just marginally heavy on the east.
Beforehand it was substantially heavy to that you could see the worm back-pedal.  No idea how this could have happened.  Very strange.


SUMMARY:  Telescope / camera ready for observing.


-          Peter (& staff).
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://mailman.tamuc.edu/pipermail/sara-reports/attachments/20161215/d3fa0a8b/attachment-0001.html>
-------------- next part --------------
A non-text attachment was scrubbed...
Name: image001.png
Type: image/png
Size: 26836 bytes
Desc: image001.png
URL: <http://mailman.tamuc.edu/pipermail/sara-reports/attachments/20161215/d3fa0a8b/attachment-0001.png>


More information about the Sara-Reports mailing list